3_4: Air from the surrounding atmosphere at 100 kPa, 20 oC, enters a compressor with a velocity of 8.6 m/s through an inlet whose diameter is 36 cm. The compressed air exits at 650 kPa, 225 oC, with a velocity of 2.8 m/s. The rate of entropy generation for the compressor is 0.062 kW/K. Determine the power input to the compressor, kW.
Answer #1

Zenit Control I volume P2 = 6sokpa 9: (0599eu = 0-060 KW / t2 = 2250 1/2 = 2.8 m/ Wsuput = ? | Toa Susalogudio up- 233:15 k 8